OpenStudy (laurenarg):

Find the coordinates for the midpoint of the segment with endpoints given. (3, 5) and (-2, 0) (1/2, 5/2) (5/2, 1/2) (1, 5/2) Could someone explain how to do this?

OpenStudy (peachpi):

Midpoint formula: \[\left( \frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 }\right)\] ^you can memorize that if you want, but basically you average the coordinates to get the midpoint

OpenStudy (laurenarg):

So the answer for this one is (5/2, 1/2)?

pooja195 (pooja195):

Not quite, lets plug in the given info

pooja195 (pooja195):

Midpoint formula: \[\left( \frac{ 3+(-2) }{ 2 },\frac{ 5+0 }{ 2 }\right)\]

pooja195 (pooja195):

Try from here; what will our answer be?

OpenStudy (laurenarg):

(1/2, 5/2)

pooja195 (pooja195):

Yep